Sandhill Co. sold $3,300,000, 7%, 10-year bonds on January 1, 2017. The bonds were dated January 1, 2017, and pay interest on January 1. The company uses straight-line amortization on bond premiums and discounts. Financial statements are prepared annually.

Prepare the journal entries to record interest expense for 2017 under both of the bond issuances assuming they sold at: (1) 104 and (2) 97
Answer #1
(1) At 104
Debit Credit
Interest expense 217800
Premium on Bonds payable 13200 =(3300000*0.04)/10
       Interest payable 231000 =3300000*7%
(2) At 97
Debit Credit
Interest expense 240900
       Discount on Bonds payable 9900 =(3300000*0.03)/10
       Interest payable 231000 =3300000*7%